OpenStudy (anonymous):

What is the x-intercept of the line perpendicular to the line defined by 3x-2y=6 and whose y-intercept is 2?

OpenStudy (anonymous):

if the line has equation 3x-2y=6 => y =3/2 x -3 then line perpendicular to above line has equation y=-2/3 x +k where k=constant=y-intercept so, y=-2/3 x + 2 now to find x-intercept, you know y = 0 so, so, x =3 i.e x-intercept = 3